OpenStudy (anonymous):

x – y = 7 x + y = 5 What is the x-value of their intersection? 6 -1 12 none of the above

OpenStudy (experimentx):

6

OpenStudy (hoblos):

x – y = 7 x + y = 5 add the two equations 2x=12 x=6

OpenStudy (anonymous):

thanks i thought it was 12

OpenStudy (anonymous):

then whats the solution

OpenStudy (anonymous):

x – y = 7 x – y = 5 What is the solution? (6, -1) no solution infinite solutions none of the above

OpenStudy (hoblos):

if you subtract the equations here you will get 0=2 impossible so no solution

OpenStudy (anonymous):

i dont get this one is it (4,2), x + 2y = 4 2x + 4y = 8 What is the solution? (4,2) no solution infinite solutions none of the above